OpenStudy (aaronandyson):

the sides of a right angled triangle are (x-1) cm . 3x cm and (3x+1)cm @hartnn

hartnn (hartnn):

and you need to find ''x" ? use pythagoras theorem, know what that is ?

OpenStudy (aaronandyson):

8 cm?

OpenStudy (aaronandyson):

I got to fins "x" the length of the sides and its area x = 8cm sides = 24cm,23cm,7cm area = 276 cm^2

OpenStudy (aaronandyson):

find*

OpenStudy (aaronandyson):

I'm right?

hartnn (hartnn):

how did you get the area? sides are correct

OpenStudy (aaronandyson):

Area = bh/2??

OpenStudy (aaronandyson):

Area = base*height/2?

hartnn (hartnn):

b and h are 23 and 7, right ?

OpenStudy (aaronandyson):

Wit.I'll draw the diagram as well 

hartnn (hartnn):

isn't it 3x+1 ?

OpenStudy (aaronandyson):

Typo sorry

hartnn (hartnn):

so 3x+1 would be the hypotenuse, because its the greatest and 3x will be the leg

OpenStudy (aaronandyson):

I calculated it again this time I got 94cm^2

hartnn (hartnn):

ooh, just realized the sides are wrong.... x= 8 so sides will be 7,24,25

OpenStudy (aaronandyson):

So,The area is 94 cm^2?

hartnn (hartnn):

Area = 24*7/2 = 84 cm^2 how you got 94? :P

OpenStudy (aaronandyson):

typos X_X

hartnn (hartnn):

LOL

OpenStudy (er.mohd.amir):

side are 25,24,7 then hypotenuse is 25 and other are legs area=(base*height)/2

hartnn (hartnn):

sometimes typos could be dangerous.. good luck in your exams!

OpenStudy (aaronandyson):

Vacation are on bro :P Refreshing my basics